OpenStudy (anonymous):

√11:(8)+√6:(12)-√5:(2) Evaluate

OpenStudy (anonymous):

\[\sqrt[11]{8}+\sqrt[6]{12}-\sqrt[5]{2}\]

OpenStudy (anonymous):

do you want a decimal approximation? or is this a middle step in another problem? (from a high school math perspective, no further simplifications are possible, because radicals of different order do not combine easily)

OpenStudy (anonymous):

i need the exact value